OpenStudy (anonymous):

how do you do this separable equations problem: dx/dt = te^t x(0)=1

OpenStudy (anonymous):

\[\frac{dx}{dt}=te^t \rightarrow dx=te^tdt \rightarrow \int\limits_{}{}dx=\int\limits_{}{}te^tdt\]

OpenStudy (anonymous):

\[\frac{dx}{dt}=te^t \rightarrow dx=te^tdt \rightarrow \int\limits_{}{}dx=\int\limits_{}{}te^tdt\]

OpenStudy (anonymous):

This is just\[x=e^t(t-1)+c\]

OpenStudy (anonymous):

To find c, you have\[x(0)=e^0(0-1)+c=1 \rightarrow c=2\]

OpenStudy (anonymous):

So your equation is\[x(t)=e^t(t-1)+2\]

OpenStudy (anonymous):

how did you get e^t(t-1) is it necessary to use the product rule for that integration or what?

OpenStudy (anonymous):

You will use integration by parts.

OpenStudy (anonymous):

could you show me how to do that?

OpenStudy (anonymous):

Set u=t and dv=e^tdt

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I'll write it and scan - it will be quicker.

OpenStudy (anonymous):

alright thanks i'm trying it now

OpenStudy (anonymous):

i'm stuck

OpenStudy (anonymous):

I'm scanning

OpenStudy (anonymous):

alright thanks

OpenStudy (anonymous):

Just have to wait a min, since the software for my scanner is a pain

OpenStudy (anonymous):

alright cool

OpenStudy (anonymous):

do you think you could help me out with another one after this?

OpenStudy (anonymous):

Maybe - I'm supposed to be getting ready for uni.

OpenStudy (anonymous):

alright well its dy/dx = (xy)^(1/2) and i just wasn't sure how you set that one up

OpenStudy (anonymous):

Just split the factors up using power laws.

OpenStudy (anonymous):

so you would end up with dy/sqrt(y) = sqrt x dx

OpenStudy (anonymous):

\[\frac{dy}{dx}=(xy)^{1/2} = x^{1/2}y^{1/2} \rightarrow \frac{dy}{y^{1/2}}=x^{1/2}dx\]

OpenStudy (anonymous):

then after you integrate would you have 2 sqrt y = 2/3 x^(3/2)

OpenStudy (anonymous):

Here's the scan for the first one.

OpenStudy (anonymous):

and when it's all said and done i had \[y= (x ^{3/2}/3)^2 + (c/2)^2\]

OpenStudy (anonymous):

i feel like thats not right at all

OpenStudy (anonymous):

\[\int\limits{}{}\frac{dy}{y^{1/2}}=\frac{y^{1/2}}{1/2}=2y^{1/2}\]

OpenStudy (anonymous):

and

OpenStudy (anonymous):

\[\int\limits_{}{}x^{1/2}dx = \frac{x^{3/2}}{3/2}=\frac{2}{3}x^{3/2}+c\]

OpenStudy (anonymous):

I added the constant then.

OpenStudy (anonymous):

Equate the two.

OpenStudy (anonymous):

\[2y^{1/2}=\frac{2}{3}x^{3/2}+c\]

OpenStudy (anonymous):

Divide by 2 to clean up:\[y^{1/2}=\frac{1}{3}x^{3/2}+c\]

OpenStudy (anonymous):

You don't have to write c/2 since c is just a constant - constants just keep absorbing constants since in the end, their value is determined by initial conditions and you'll end up with the same number.

OpenStudy (anonymous):

ohhhh ok i think i got it since c is a constant i dont need to divide it

OpenStudy (anonymous):

haha ya thanks alot

OpenStudy (anonymous):

np

OpenStudy (anonymous):

i really appreciate it

OpenStudy (anonymous):

become a fan then :)

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

there should be a link next to my name in this area we're typing in. It's a blue link that says, "Become a fan"

OpenStudy (anonymous):

Should have a 'thumbs up' icon...

OpenStudy (anonymous):

ya i dont see it?

OpenStudy (anonymous):

it just has hero and the 110 fans

OpenStudy (anonymous):

refresh the page...sometimes that helps

OpenStudy (anonymous):

it just has hero and the 110 fans

OpenStudy (anonymous):

yeah, the site's awkward

OpenStudy (anonymous):

ah

OpenStudy (anonymous):

sweet i got it

OpenStudy (anonymous):

happy travels )

OpenStudy (anonymous):

thanks man i appreciate it

OpenStudy (anonymous):

later